# Stop Overthinking: A Survey on Efficient Reasoning for Large Language Models

### v1

@@ 1 Introduction
Large Language Models (LLMs) have emerged as exceptionally powerful AI tools, demonstrating advanced capabilities in natural language understanding and complex reasoning. Recently, the emergence of reasoning-focused LLMs, also referred to as Large Reasoning Models (LRMs) [91], such as OpenAI o1 [61] and DeepSeek-R1 [31], has significantly enhanced their performance in System-2 reasoning domains [44, 8], including mathematics [16, 35] and programming [17, 7]. Evolving from foundational pre-training models (e.g., LLaMA [80, 30], Qwen [95]) with next-token prediction training [23], these models leverage Chain-of-Thought (CoT) [86] prompting to generate explicit, step-by-step reasoning sequences before arriving at a final answer, considerably improving their effectiveness in reasoning-intensive tasks.
Reasoning abilities in LLMs are typically developed through supervised fine-tuning (SFT) and reinforcement learning (RL), which promote iterative and systematic problem-solving abilities. Specifically, the OpenAI o1 [61] training pipeline potentially [^1] Based on public speculation since GPT-o1 has not released detailed training methods. incorporates SFT and RL with Monte Carlo Tree Search (MCTS) [71] and a processed reward model (PRM) [47]. DeepSeek-R1 is initially fine-tuned using SFT with Long CoT reasoning data generated by RL-trained DeepSeek-R1-Zero, and then further refined through RL with a rule-based reward function.
@@
However, while long CoT reasoning significantly enhances reasoning capabilities and accuracy, the introduction of the CoT-like mechanisms (e.g., Self-consistency [84], Tree-of-Thought [96], Incentivizing RL [31]) also results in lengthy output responses, leading to substantial computational overhead and thinking time. For instance, “overthinking problem” happens when asking “what is the answer of 2 plus 3?” [10] to OpenAI-o1, DeepSeek-R1, and QwQ-32B-Preview. The reasoning sequences from such models can sometimes span thousands of tokens, many of which are redundant and do not substantially contribute to deriving the correct answer. Such verbosity directly raises inference costs and latency, restricting the practical use of reasoning models in computation-sensitive real-world applications, including real-time autonomous driving systems, interactive assistants, robotics control, and online search engines.
@@
Efficient reasoning, particularly the reduction of reasoning length, offers significant benefits, such as cutting costs and enhancing reasoning capabilities in real-world deployments. Recently, numerous studies [54, 98, 32, 56, 33] have attempted to develop more concise reasoning paths, making efficient reasoning a prominent and rapidly evolving area of research.
@@
In this paper, we provide the first structured survey to systematically explore the current progress of efficient reasoning for LLMs. As shown in Figure 2, we categorize existing works into below key directions: (1) model-based efficient reasoning, which involves optimizing full-length reasoning models into more concise reasoning models or directly fine-tuning models to achieve efficient reasoning; (2) reasoning output-based efficient reasoning, which dynamically reduces reasoning steps and output length during inference; (3) input prompts-based efficient reasoning, which improves reasoning efficiency by leveraging prompt properties, such as prompt-guided length or prompt difficulty. Orthogonal to the definition of model compression techniques in LLMs, such as quantization [27, 48] or kv cache compression [103, 52], which focus on compressing the model size and enabling lightweight inference, efficient reasoning in LLMs emphasizes Smart and Concise Reasoning by optimizing the reasoning length and reducing thinking steps. Overall, we summarize the efficient reasoning methods into below categories:
@@
  • RL with Length Reward Design (Section 3.1);
@@
  • SFT with Variable-Length CoT Data (Section 3.2);
@@
  • Compressing Reasoning Steps into Fewer Latent Representation (Section 4.1);
  • Dynamic Reasoning Paradigm during Inference (Section 4.2);
  • Prompt-guided Efficient Reasoning (Section 5.1);
  • Routing Prompts to Optimize Reasoning Efficiency (Section 5.2);
Additionally, in this paper, we explore the other interesting topics, including
  • Training Reasoning Models with Efficient Data (Section 6.1);
@@
  • Reasoning Ability of Small Language Models and Model Compression (Section 6.2);
  • Evaluation and Benchmark of Efficient Reasoning Models (Section 7);
We will continuously update our public repository with the latest research on efficient reasoning.
@@
, ERTleaf node] ] ] [ Reasoning Output-based Efficient Reasoning , OER node [ Latent Representation Compression , OER node [e.g. Coconut [33]; CODI [70]; CCoT [12]; Heima [68]; Token Assorted [74]; Loop [66]; SoftCoT [93]
, OERleaf node] ] [ Dynamic Reasoning Paradigm , OER node [e.g. Speculative Rejection [76]; Sampling-Efficient TTS [85]; DPTS [24]; Certaindex [28]; Fast MCTS [41]; ST-BoN [85]; More is Less [88]; RSD [45]; LightThinker [101]; INFTYTHINK [94]; SCoT [90]; RASC [83]; Adaptive Reasoning [100]; AdaptiveStep [51]; , OERleaf node] ] ] [ Input Prompts-based Efficient Reasoning , ERI node [ Prompt-Guided Efficient Reasoning , ERI node [e.g. Token-Budget [32]; Chain of Draft [92]; Token Complexity [40]; Concise Chain-of-Thought (CCoT) [65] , ERIleaf node] ] [ Routing by Question Attributes , ERI node [e.g. Claude 3.7 Sonnet [2]; SoT [4]; , ERIleaf node] ] ] [ Efficient Data and Models , CP node [ Less Training Data , CP node [e.g. LIMO [97]; s1 [58]; S2[55]; Light-R1 [87]; , CPleaf node] ] [ Pruning & Quantization & Distillation , CP node [e.g. Struggle[43]; Strong Verifiers [73]; TinyR1-32B-Preview [77]; Mixed Distillation [13]; Counterfactual Distillation [26]; Feedback-Driven Distillation [105]; SKIntern [46]; Adaptive Thinking [9]; PRR [104]; , CPleaf node] ] ] [ Benchmark & Insights , BI node [ Evaluation & Benchmarks , BI node [e.g. 1B vs. 405B [49]; Sys2Bench [62]; Danger [19]; , BIleaf node] ] ] ]
@@
2 Background: Long CoT Reasoning Models and Overthinking Phenomenon
2.1 Chain-of-Thought (CoT)
CoT [86] is one key approach by which reasoning capabilities have been purposely introduced in LLMs. In this setting, models are often prompted to output a reasoning thinking chain before arriving at a final answer. Techniques proposed in this realm have been shown to improve overall accuracy-like performance [86] since via improving the quality of the generation context, a more consistent final result is likely to be achieved.
Some notable CoT variants include Self-consistency CoT [84], which replaces the greedy-like thinking chain generation by sampling more diverse reasoning paths and then selects the most consistent answer by marginalizing and aggregating such paths. Similarly, Tree-of-Thought prompting [96] proposes problem-solving by structuring thought processes into a tree with backtracking allowed, which greatly improves the efficiency of parallelizable subtasks. Graph-of-Thoughts [6] prompting further extends this concept by structuring thoughts as a graph and permits looping to refine an individual thought. There exist many further variants of CoT, but generally, they are done within the scope of prompting the model in different ways to enable different behaviors and (optionally) implementing a controller-like component to manage the progression of thought generation and usage.
@@ 2.2 The Mechanism behind Large Reasoning Models
Multi-step reasoning capability refers to the ability of LLMs to actively generate a reasoning process before committing a final answer. Reasoning steps can significantly improve LLM inference, especially in logic-intensive tasks such as math and programming. On a broader scale, reasoning-capable model LLMs are typically favored by human users than their non-reasoning-capable counterparts, as showcased in the Chatbot Arena LLM Leader board [^2] A community-driven evaluation for best LLM and AI chatbots https://lmarena.ai/?leaderboard.
@@
Recent reasoning models, such as DeepSeek-R1 [31] and OpenAI o1 [54], are known (or speculated) to internally learn reasoning behaviors without relying on explicit and computationally intensive test-time argumentations. These models can produce detailed CoT reasoning by iteratively generating intermediate steps and progressively refining solutions through sequentially decoding until reaching a final answer. This approach is fundamentally different from the prompting-based strategy widely adopted in CoT variants, as the reasoning capability is “internalized” within those reasoning-capable models via training.
@@
OpenAI o1 model, as speculated, leverages a tree-based approach such as MCTS [38, 18] with a Process Reward Model to be trained to systematically explore various reasoning paths and determine the optimal solutions through guided simulations [^3] We emphasize that there is no official confirmation regarding the training details and mechanism of OpenAI o1, so we are merely presenting one popular speculation here. Sources like and discuss this topic in-depth and are recommended to interested readers. Models like DeepSeek-R1 are specifically trained to enable its reasoning capability. Through a combination of supervised fine-tuning and reinforcement learning recipes with a special focus on rule-based rewards upon math and coding tasks, DeepSeek-R1 and the like learn to generate the reasoning process within a predefined format before providing the final answer.
2.3 Overthinking Problem in Long CoT Reasoning Models
@@
Figure 4: An example of the “overthinking phenomenon”: when the reasoning model is asked “Which is larger, 0.9 or 0.11?”, it takes an unnecessarily long time (i.e., 19 seconds for QwQ-32B [79] and 42 seconds for DeepSeek-R1 [31]) to deliver its final answer. The example was tested in March 2025.
@@
“Overthinking phenomenon” [78, 10] in long CoT reasoning refers to the phenomenon in which LLMs generate overly detailed or unnecessarily elaborate reasoning steps, thereby impairing their problem-solving efficiency. Specifically, some modern reasoning LLMs — especially some reasoning-capable offerings at smaller parameter scales — tend to produce verbose reasoning or redundant intermediate steps, making them incapable of providing an answer before the user-defined token budget runs out or worse: providing an incorrect answer because the verbose reasoning steps introduce errors or diminishes the logical clarity.
Figure 4 showcases the example of overthinking. Although overthinking CoT reasoning produces the correct answer within its initial steps, its generated reasoning steps often include multiple redundant intermediate steps, resulting in unnecessary complexity and reduced clarity. Given the extreme resource cost associated with the decoding phase of LLMs (e.g., OpenAI o1 costs \$60 per 1M of generated tokens), this is highly undesirable behavior; and it goes without saying that it would be even worse if an extended reasoning generation leads to a wrong answer. In contrast, efficient thinking with fewer steps can gain the correct answers, which highlights the possibility of reducing the extra budget of overthinking tokens.
Note that this challenge of efficient reasoning is considered significant because the pretraining recipes of reasoning-capable models often explicitly encourage models to generate extended reasoning steps to pursue the correct answer. For instance, when DeepSeek-R1-Zero is trained longer, its response length and benchmark performance both grow [31]; where observing such two trends are often considered the proxy of a successful reasoning-enabling training recipe. Thus, wanting to achieve reasoning efficiency at inference is, by design, working against certain pre-training goals of the model and thereby requires non-trivial considerations. This work aims to summarize the different schools of thought, as well as their exemplary methods, in achieving the challenging but rewarding goal of having an efficient yet capable reasoning model.
@@ 3.1 RL with Length Reward Design
KIMI [78] PPO $\begin{cases}\;\lambda&\!\text{if correct},\lambda=0.5\!-\!\frac{L(y_{\text{pred}})-L_{\min}}{L_{\max}-L_{\min}}\\ \;\min(0,\lambda)&\!\text{if wrong}.\end{cases}$
L1 [1] PPO $\begin{cases}\;x_{\text{new}}=\textit{CONCAT}~{}(x,\textit{\textquotedblleft Thinkfor N tokens.\textquotedblright}),\\ \;r(y,y_{GT},L(y_{GT}))=\mathbb{I}(y_{\text{pred}}=y_{GT})-\alpha\cdot|L(y_{GT})-L(y_{\text{pred}})|\end{cases}$ AMC GPQA LAST MMLU MATH-500 AIME-2024 Olympiad-Bench DeepSeek-R1-Distill-Qwen-1.5B
@@
Table 2: Comparison of different policy optimization methods in CoT length controls. $\hat{R}_{t}$ represents the reward model. $\pi_{\text{ref}}$ is the policy of reference model. $\gamma$ is a target reward margin term for SimPO. The $y_{w}$ is for winning responses and $y_{l}$ is for losing responses.
@@
Existing works leverage traditional RL optimization techniques combined with explicit length-based reward to control the length of CoT reasoning. Some detailed length rewards are shown in Table 1. The work [3] proposes utilizing length-based rewards conditioned on correctness, where shorter correct answers receive higher rewards. They then apply traditional policy gradient methods guided by this reward scheme to encourage LLMs to produce concise reasoning steps. Expanding from the policy gradient, the following discussed work is primarily built upon proximal policy optimization (PPO) [67] with CoT length penalty. Demystifying [98] presents empirical findings from RL experiments examining how reasoning capability is influenced by length. They demonstrate that RL does not consistently or reliably increase the length and complexity of CoT reasoning, emphasizing the necessity of controlling CoT length growth to ensure stable performance. To mitigate these issues, they proposed a Cosine Reward based on a Dirichlet function of concise reward formula [53] and the proposed “exceed length penalty” scores. Due to the performance impact of CoT length, Kimi k1.5 [78] incorporates length penalty in the policy optimization for long-CoT activations and model merging. Besides optimizing with length penalty reward, L1 [1] modify the training data with the designated length constraint instruction (i.e., Think for $N$ tokens) before launching the policy optimization with pre-trained reasoning LLMs. O1-Pruner [54] introduces the Length-Harmonizing Reward, combined with a PPO-style loss, to optimize reasoning LLMs by effectively shortening the CoT length. Specifically, the Length-Harmonizing Reward is computed based on the ratio of CoT lengths between the reference model output and the predicted results. Additionally, this reward incorporates accuracy-based constraints comparing predictions to the reference model outputs, ensuring that shortening the reasoning process does not degrade task performance. Without relying on a reference model, DAST [69] employs SimPO [57] to fine-tune reasoning LLMs using a constructed length-preference dataset. This dataset is generated based on a self-defined token-length budget measurement $L_{\text{budget}}$, defined as a linear combination of the average token length of correct responses and the maximum allowed generation length.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have emerged as exceptionally powerful AI tools, demonstrating advanced capabilities in natural language understanding and complex reasoning. Recently, the rise of reasoning-focused LLMs, also referred to as reasoning-capable models or Large Reasoning Models (LRMs) [98] such as OpenAI o1 [65] and DeepSeek-R1 [33], has significantly improved performance in System-2 reasoning domains [47], particularly in challenging mathematics [18, 37] and programming tasks [19, 7]. Evolving from foundational pretrained models (e.g., LLaMA [86, 32]) trained with next-token prediction [25], these models typically leverage Chain-of-Thought (CoT) [92] reasoning chains to generate explicit, step-by-step reasoning sequences before arriving at a final answer, significantly improving their effectiveness in reasoning-intensive tasks.
Such reasoning abilities in LLMs are typically developed through supervised fine-tuning (SFT) and reinforcement learning (RL), which promote iterative and systematic problem-solving abilities. For instance, DeepSeek-R1 [33] undergoes multiple rounds of SFT and RL training, emphasizing structured thinking templates and rule-based reward mechanisms. In particular, the rule-based rewards provides precise and explicit feedback signals during training, effectively enhancing the general reasoning capabilities beyond the pretrained LLM.
@@
However, while long CoT reasoning significantly boosts accuracy, step-by-step thinking mechanisms also lead to lengthy output responses, resulting in substantial computational overhead and increased reasoning time. For instance, the "overthinking problem" arises when answering a simple question [10] like, "what is the answer of 2 plus 3?" Some reasoning models, especially smaller ones, can generate reasoning sequences spanning thousands of tokens. This verbosity significantly increases both inference costs and latency, limiting the practical application of reasoning models in computation-sensitive real-world scenarios, such as real-time autonomous driving systems, interactive conversational assistants, precision robotic control tasks, and large-scale online search engines.
@@
Efficient reasoning, particularly the reduction of reasoning length, offers significant benefits in such regards, providing direct cost reduction and improved feasibility for real-world deployments. Recently, numerous studies [57, 104, 34, 59, 35] have explored ways to develop more concise reasoning paths, making efficient reasoning a rapidly evolving research area.
@@
In this paper, we present the first structured survey systematically exploring the progress in efficient reasoning for LLMs. As illustrated in Figure 2, we categorize existing work into three key directions: (1) Model-based efficient reasoning, which focuses on optimizing full-length reasoning models into more concise variants or directly training efficient reasoning models. (2) Reasoning output-based efficient reasoning, which dynamically reduces reasoning steps and length during inference. (3) Input prompts-based efficient reasoning, which enhances reasoning efficiency based on input properties such as difficulty or length control. Unlike model compression techniques such as quantization [97, 29, 50] or KV cache compression [110, 55, 76, 107], which focus on reducing model size for lightweight inference, efficient reasoning in LLMs emphasizes smart and concise reasoning by optimizing the length of generated reasoning sequences and reducing unnecessary thinking steps.
Overall, we provide a summary of the current key approaches to efficient reasoning, organizing them into following categories:
@@
  • Reinforcement Learning with Length-Based Reward Design (Section 3.1)
@@
  • Supervised Fine-Tuning with Variable-Length CoT Data (Section 3.2)
@@
  • Compressing Reasoning Steps into Fewer Latent Representations (Section 4.1)
  • Dynamic Reasoning Paradigms During Inference (Section 4.2)
  • Prompt-Guided Efficient Reasoning (Section 5.1)
  • Routing Prompts to Optimize Reasoning Efficiency (Section 5.2)
Additionally, we explore other relevant topics, including:
  • Training Reasoning Models with Efficient Data (Section 6.1)
@@
  • Reasoning Abilities of Small Language Models and Model Compression (Section 6.2)
  • Evaluation and Benchmarking of Efficient Reasoning Models (Section 7)
@@
, ERTleaf node] ] ] [ Reasoning Output-based Efficient Reasoning , OER node [ Latent Representation Compression , OER node [e.g. Coconut [35]; CODI [75]; CCoT [12]; Heima [73]; Token Assorted [79]; Loop [70]; SoftCoT [100];
, OERleaf node] ] [ Dynamic Reasoning Paradigm , OER node [e.g. Speculative Rejection [81]; Sampling-Efficient TTS [91]; DPTS [26]; Certaindex [30]; Fast MCTS [44]; ST-BoN [91]; More is Less [94]; RSD [48]; LightThinker [108]; INFTYTHINK [101]; SCoT [96]; RASC [89]; Adaptive Reasoning [106]; AdaptiveStep [54]; , OERleaf node] ] ] [ Input Prompts-based Efficient Reasoning , ERI node [ Prompt-Guided Efficient Reasoning , ERI node [e.g. Token-Budget [34]; Chain of Draft [99]; Token Complexity [43]; Concise Chain-of-Thought (CCoT) [69]; MARP [8]; , ERIleaf node] ] [ Routing by Question Attributes , ERI node [e.g. Claude 3.7 Sonnet [2]; SoT [4]; Self-REF [17]; Confident [14]; RouteLLM [64]; , ERIleaf node] ] ] [ Efficient Data and Models , CP node [ Less Training Data , CP node [e.g. LIMO [103]; s1 [61]; S2[58]; Light-R1 [93]; , CPleaf node] ] [ Pruning & Quantization & Distillation , CP node [e.g. Struggle[46]; Strong Verifiers [78]; TinyR1-32B-Preview [82]; Mixed Distillation [13]; Counterfactual Distillation [28]; Feedback-Driven Distillation [112]; SKIntern [49]; Adaptive Thinking [9]; PRR [111]; , CPleaf node] ] ] [ Benchmark & Insights , BI node [ Evaluation & Benchmarks , BI node [e.g. 1B vs. 405B [52]; Sys2Bench [66]; Danger [21]; Inference-time Computation [51]; Impact [39]; , BIleaf node] ] ] ]
@@
2 Background: Long CoT Reasoning Models and the Overthinking Phenomenon
2.1 Chain-of-Thought (CoT) Reasoning
Chain-of-Thought (CoT) reasoning [92] is a key approach that has been purposefully introduced in LLMs to enhance their reasoning capabilities. In this setting, models are typically prompted to generate a structured reasoning chain before arriving at a final answer. Techniques in this domain have been shown to improve overall accuracy [92] since a higher-quality generation context often leads to more consistent and reliable final results. Several notable CoT variants have been developed: Self-Consistency CoT [90] replaces the standard greedy decoding approach by sampling diverse reasoning paths and selecting the most consistent answer through marginalization and aggregation. Tree-of-Thought (ToT) prompting [102] further structures the reasoning process as a tree with backtracking, significantly improving efficiency in solving parallelizable subtasks. Graph-of-Thoughts (GoT) prompting [6] extends this concept by structuring thoughts into a graph, allowing iterative refinement of individual reasoning steps. While many CoT variants exist, they generally involve different prompting techniques to guide the behavior of models, sometimes incorporating controller-like mechanisms to manage thought progression and usage.
@@ 2.2 The Mechanism Behind Large Reasoning Models
Multi-step reasoning refers to the ability of LLMs ability to generate structured reasoning steps before committing to a final answer. This capability is particularly beneficial for logic-intensive tasks such as mathematics and programming. More broadly, reasoning-capable models are often favored by human users over their non-reasoning counterparts, as evidenced by rankings in the Chatbot Arena LLM Leaderboard. [^1] A community-driven evaluation of leading LLMs and AI chatbots: https://lmarena.ai/?leaderboard.
@@
Recent reasoning models, such as DeepSeek-R1 [33] and OpenAI o1 [57], are known or believed to have internalized reasoning behaviors, reducing reliance on explicit test-time augmentations. These models generate detailed CoT reasoning by iteratively producing intermediate steps and refining solutions sequentially until reaching a final answer. Unlike traditional CoT approaches, which rely on prompting, these reasoning models internalize their reasoning capability through extensive training.
@@
The OpenAI o1 model is speculated to employ a tree-based search approach, such as Monte Carlo Tree Search (MCTS) [41, 20], combined with a Process Reward Model (PRM) to explore reasoning paths and determine optimal solutions through guided simulations. [^2] There is no official confirmation regarding OpenAI o1’s training details and mechanisms. However, sources such as https://www.interconnects.ai/p/openais-o1-using-search-was-a-psyop and https://www.youtube.com/watch?v=6PEJ96k1kiw discuss these speculations in detail and are recommended for interested readers. DeepSeek-R1, on the other hand, explicitly learns its reasoning capability through supervised fine-tuning and reinforcement learning, with a particular emphasis on rule-based rewards for math and coding tasks. These models are trained to generate reasoning steps in a predefined format before arriving at their final answers.
2.3 The Overthinking Problem in Long CoT Reasoning Models
@@
Figure 4: An example of the “overthinking phenomenon”: when asked “Which is larger, 0.9 or 0.11?”, the reasoning model takes an unnecessarily long time (e.g., 19 seconds for QwQ-32B [84] and 42 seconds for DeepSeek-R1 [33]) to arrive at the correct answer. This example was tested in March 2025.
@@
The “overthinking phenomenon” [83, 10] in long CoT reasoning models refers to situations where LLMs generate excessively detailed or unnecessarily elaborate reasoning steps, ultimately reducing their problem-solving efficiency. In particular, many modern reasoning models, especially those with smaller parameter scales, tend to produce verbose reasoning or redundant intermediate steps, making them unable to provide answers within the user-defined token budget. In worse cases, excessive reasoning steps introduce errors or obscure logical clarity, leading to incorrect answers.
Figure 4 illustrates an example of overthinking. Even though the model arrives at the correct answer early in its reasoning process, it continues generating unnecessary intermediate steps, leading to inefficiencies. Given the substantial resource costs associated with LLM inference (e.g., OpenAI o1 costs \$60 per 1M generated tokens), such behavior is highly undesirable. Moreover, the problem becomes even worse if longer reasoning leads to wrong answers. In contrast, efficient reasoning models would use fewer reasoning steps to obtain correct answers while reducing inference costs.
Addressing this challenge is particularly difficult because the pretraining recipes for reasoning-capable models often explicitly encourage generating extended reasoning steps to improve accuracy. For example, DeepSeek-R1-Zero, a more or less a development prototype of DeepSeek-R1, exhibits a direct correlation between increased training duration with longer response lengths and improved benchmark performance [33]. These trends are often viewed as proxies for successful reasoning training. Consequently, improving inference efficiency requires working against certain pretraining objectives, making it a non-trivial challenge.
This paper aims to systematically summarize various approaches and methodologies toward achieving the challenging yet valuable goal of developing reasoning models with high efficiency and strong reasoning capabilities.
@@ 3.1 RL with Length Reward Design
L1 [1] GRPO $\begin{cases}\;x_{\text{new}}=\textit{CONCAT}~{}(x,\textit{\textquotedblleft Thinkfor N tokens.\textquotedblright}),\\ \;r(y,y_{GT},L(y_{GT}))=\mathbb{I}(y_{\text{pred}}=y_{GT})-\alpha\cdot|L(y_{GT})-L(y_{\text{pred}})|\end{cases}$ AMC GPQA LAST MMLU MATH-500 AIME-2024 Olympiad-Bench DeepSeek-R1-Distill-Qwen-1.5B
@@
Table 2: Comparison of different policy optimization methods in CoT length controls. $\hat{R}_{t}$ represents the reward model. $\pi_{\text{ref}}$ is the policy of reference model. $\gamma$ is a target reward margin term for SimPO. The $y_{w}$ is for winning responses and $y_{l}$ is for losing responses, where some with $G$ on superscript denote the outputs of different sampled groups.
@@
GRPO [72] $\mathbb{E}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y^{G}_{w}\mid x)}{\pi_{\text{ref}}(y^{G}_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y^{G}_{l}\mid x)}{\pi_{\text{ref}}(y^{G}_{l}\mid x)}\right)-\lambda\mathbb{D}_{KL}[\pi_{\theta}||\pi_{\text{ref}}]\right]$
@@
Existing works leverage traditional RL optimization techniques combined with explicit length-based reward to control the length of CoT reasoning. Some detailed length rewards are shown in Table 1. The work [3] proposes utilizing length-based rewards conditioned on correctness, where shorter correct answers receive higher rewards. They then apply traditional policy gradient methods guided by this reward scheme to encourage LLMs to produce concise reasoning steps. Expanding from the policy gradient, the following discussed work is primarily built upon proximal policy optimization (PPO) [71] with CoT length penalty. Demystifying [104] presents empirical findings from RL experiments examining how reasoning capability is influenced by length. They demonstrate that RL does not consistently or reliably increase the length and complexity of CoT reasoning, emphasizing the necessity of controlling CoT length growth to ensure stable performance. To mitigate these issues, they proposed a Cosine Reward based on a Dirichlet function of concise reward formula [56] and the proposed “exceed length penalty” scores. Due to the performance impact of CoT length, Kimi k1.5 [83] incorporates a length penalty into its policy optimization (a variant of online policy mirror decent [85]) to improve long CoT activations and facilitate effective model merging. Besides optimizing with length penalty reward, L1 [1] modify the training data with the designated length constraint instruction (i.e., Think for $N$ tokens) before launching the policy optimization with pre-trained reasoning LLMs. O1-Pruner [57] introduces the Length-Harmonizing Reward, combined with a PPO-style loss, to optimize reasoning LLMs by effectively shortening the CoT length. Specifically, the Length-Harmonizing Reward is computed based on the ratio of CoT lengths between the reference model output and the predicted results. Additionally, this reward incorporates accuracy-based constraints comparing predictions to the reference model outputs, ensuring that shortening the reasoning process does not degrade task performance. Without relying on a reference model, DAST [74] employs SimPO [60] to fine-tune reasoning LLMs using a constructed length-preference dataset. This dataset is generated based on a self-defined token-length budget measurement $L_{\text{budget}}$, defined as a linear combination of the average token length of correct responses and the maximum allowed generation length.